OpenStudy (courtney):

what is the equationof a line with a slope -3/4 passing through the point (-2,-10)

OpenStudy (anonymous):

The equation of a line is y=mx+b. You are given an x, a y and an m. You can then solve for b. Then write the equation as y=(m you are given)x+(b you solved for)

OpenStudy (courtney):

i dont get math at all and you just confused me a whole lot

OpenStudy (anonymous):

Sorry. The equation of any like is y=mx+b where m is the slope and b is the y-intercept. They give you the slope which is m and the coordinates (-2,-10) are in the form (x,y). That means they gave you the values for m, x and y. So now you have -10=(-3/4)(-2)+b. You then solve for b. Then write the equation as y=(m which is the slope)x+(b you solved for)

OpenStudy (anonymous):

line*

OpenStudy (courtney):

ok can you explain the very last part again

OpenStudy (anonymous):

Sure. So now you solved for b and you are given m. y and x remain as variables though. So you write y=mx+b but plug in the m they gave you and the b you solved for. The x and y remain as just x and y.

OpenStudy (courtney):

ok i still dont get it could you just give me the answer please lol

OpenStudy (anonymous):

I really want you to get it =/ -10=(-3/4)(-2)+b. You then solve for b. y=mx+b y=(-3/4)x-(23/2) Hope you see how i got to that

OpenStudy (courtney):

no but could you help me with something else

OpenStudy (anonymous):

At what step do you get lost.

OpenStudy (courtney):

basically the whole thing sorry

OpenStudy (anonymous):

So you know that the equation of any line is y=mx+b. Where m is the slope of the line and b is the point where it crosses the y axis, right.

OpenStudy (courtney):

yea

OpenStudy (anonymous):

So in the problem they give you the slope and the coordinate of a point that is supposed to be on this line. So they give you m a certain x and a certain y

OpenStudy (courtney):

uhh ok im getting lost

OpenStudy (anonymous):

"what is the equation of a line with a slope -3/4 passing through the point (-2,-10)" Slope is -3/4. Since slope is the m value we have y=(-3/4)x+b, right.

OpenStudy (courtney):

im really bad at math so sorry if im wasting your time

OpenStudy (anonymous):

Don't worry about it. Do you get what i did on the last step though? we had y=mx+b as the equation of the line. They told us the slope which is m so I just plugged it in.

OpenStudy (courtney):

ok sort of

OpenStudy (anonymous):

So we have y=(-3/4)x+b. Now they tell us that this point (-2,-10) is on the line. The point is written in the form (x,y). Ok so far?

OpenStudy (courtney):

ok u can keep goin

OpenStudy (anonymous):

So we know that since this point is on the line, it makes the equation of the line true. This means that -10=m(-2)+b. Since we are also given m, we have -10=(-3/4)(-2)+b.

OpenStudy (courtney):

ok im starting to get it

OpenStudy (anonymous):

Now we can solve for b since everything else is just numbers. -10=(3/2)+b. -10-(3/2)=b b=-23/2 So now we have m=-3/4 and b=-23/2

OpenStudy (courtney):

ok

OpenStudy (anonymous):

Now since the equation of the line is just y=mx+b where y and x are variables, we have all the information we need. Now we just plug them in: y=(-3/4)x+(-23/2) y=(-3/4)x-(23/2)

OpenStudy (courtney):

ok is that all

OpenStudy (anonymous):

Yup. You now have the equation of the line that has a slope -3/4 passing through the point (-2,-10)

OpenStudy (courtney):

omg thank u so much are you an adult

OpenStudy (anonymous):

No problem Almost :P. Still in high school

OpenStudy (courtney):

oh ill be in highschool this coming year

OpenStudy (courtney):

are you on facebook